# Paragraph-to-Image Generation with Information-Enriched Diffusion Model

### v1

@@ 4.2.3 ParaPrompts-400
w LoRA 4.2M (0.06%) 50.6 (+5.5) 4.3 45.1
w LoRA 16.7M (0.25%) 51.1 (+5.8) 3.6 45.3
w LoRA 33.6M (0.51%) 53.4 (+4.7) 2.1 48.7
w LoRA 67.1M (1.01%) 51.9 (+6.6) 2.8 45.3
Table 4: Ablation Study for ParaImage on ParaPrompts-400.
@@
w/o ParaImage (Pre-Train, Stage-1) - - - -
w ParaImage-B (Stage-2 $vs.$ Stage-1) 82.1 17.9 76.9 32.1
w ParaImage-S (Stage-3 $vs.$ Stage-2) 86.2 13.8 53.1 46.9
@@ 4.3 Ablation Study
4.3.1 Effect of Language Model Adaptation
@@
Table 3 presents the ablation study for LLM Adaptation. ‘Base’ denotes directly performing paragraph-image alignment learning (Stage-2) with ParaImage-Big without fine-tuning Llama V2 using LoRA. Compared to the base model, our LLM adaptation with LoRA demonstrates an improvement of nearly $5\%$ in human voting rates. Another insight reveals that during the process of increasing trainable parameters from $4.2$ million to $67.1$ million, the performance appears to gain consistently. Therefore, in our final configuration, we randomly select 16.7 million trainable parameters as the ultimate setting. Figure 9 (b) presents the visual comparisons for the performance from LLM adaptation.
4.3.2 Effect of ParaImage
@@
Table 4 presents the ablation study for the proposed dataset, ParaImage. ParaImage-Big brings significant improvements, with approximately a $50\%$ increase in human voting rates for both visual appeal and text faithfulness simultaneously. The smaller-scale, and high-quality set of 3k images in ParaImage-Small further enhances aesthetic aspects, achieving a remarkable around $70\%$ increase in human preference rates compared to that of Stage 2. Figure 9 (a) presents a visual comparison, vividly illustrating the gains in visual appeal. In this case, the prompt is ‘An exquisite sculpture in the center of a square on a rainy day.’
@@ 6.1 More Visualizations
Table 10 and Table 11 provides more visualizations of ParaDiffusion for human-centric and scenery-centric domains, respectively. The visualization reveals that our ParaDiffusion can generate intricate and realistic composite images of individuals as well as picturesque scenes. Moreover, it is noteworthy that images generated through paragraph-image generation exhibit a compelling narrative quality, enriched with profound semantics. The generated images consistently feature intricate object details and demonstrate effective multi-object control.
@@ 6.3 Experiments on 1,600 prompts of PartiPrompts
We also provides the related experiment results on PartiPrompts-1600, as shown in Figure 13. It can be observed that our model also achieved outstanding performance in Text Faithfulness, with a $27.3\%$ human voting rate, significantly outperforming previous models such as SD XL and DeepFloyd-IF. Additionally, our model demonstrated a competitive advantage in visual appeal, surpassing SD XL and DeepFloyd-IF, and approaching the performance of contemporaneous work PIXART-$\alpha$. A notable observation in ParaPrompts is the high proportion of ’Tie’ votes in human evaluations, especially for Text Faithfulness, with a voting rate of up to $34\%$. This is attributed to the presence of numerous simple or abstract prompts in ParaPrompts-1600, making it challenging to provide precise voting results, such as for prompts like ‘happiness’ or ‘emotion.’

### v3

@@ 4.2 Performance Comparisons and Analysis
w LoRA ($r$=8, $\alpha$=32) 4.2M (0.06%) 50.6% 4.3% 45.1%
w LoRA ($r$=16, $\alpha$=32) 8.4M (0.13%) 50.8% 2.6% 46.6%
w LoRA ($r$=32, $\alpha$=32) 16.7M (0.25%) 51.1% 3.6% 45.3%
w LoRA ($r$=64, $\alpha$=32) 33.6M (0.51%) 53.4% 2.1% 44.5%
w LoRA ($r$=128, $\alpha$=32) 67.1M (1.01%) 51.9% 2.8% 45.3%
w LoRA ($r$=64, $\alpha$=16) 33.6M (0.51%) 52.3% 2.5% 45.2%
w LoRA ($r$=64, $\alpha$=32) 33.6M (0.51%) 53.4% 2.1% 44.5%
w LoRA ($r$=64, $\alpha$=64) 33.6M (0.51%) 53.1% 2.8% 44.1%
Table 7: Definition and explanation of LoRA Hyperparameters $r$ and $\alpha$.
Params Definition Effect
$r$ (LoRA Rank) Dimension of low-rank decomposition Higher $r$ improves expressivity but increases parameters and computation.
$\alpha$ (LoRA Alpha) Scaling factor for LoRA updates Higher $\alpha$ strengthens LoRA’s impact but may affect stability.
$\frac{\alpha}{r}$ (Scaling) Scaling applied to $AB$ update Adjusts the update magnitude, balancing adaptation and stability.
Table 8: Ablation for Image Quality on ParaPrompts-400. $3$k training data is used for all datasets. We only evaluate Visual Appeal.
Dataset Voting
LAION 3.1%
LAION-Aesthetics 18.5%
ParaImage-Big 23.7%
ParaImage-Small 54.7%
Table 9: Ablation for Data Size of ParaImage-Big on ParaPrompts-400 (Human Voting) and MS-COCO $256\times 256$ (FID-30K). We only evaluate Visual Appeal for Human Voting. We randomly sample the data to a smaller scale.
[…]
0.3m 17.1% 13.45
1.3m 32.7% 10.23
3.3m 50.2% 9.64
4.2.4 Layout-based Image Generative v.s. Text-based Image Generation
The current approaches to paragraph-to-image generation can be broadly categorized into two main strategies. The first is layout-based image generation, which utilizes models like ChatGPT or GPT-4 to generate bounding box layouts as the condition for image generation, as illustrated in Figure 10. The second approach is our pure text-to-image alignment method for paragraph-to-image generation.
To better compare and analyze these two paradigms, we provide corresponding experimental results and visualizations, as shown in Figure 9 and Figure 11. LLM Blueprint [8], the SOTA layout-based model, as a baseline for comparison with our proposed model. Figure 9 demonstrates that our model significantly outperforms LLM Blueprint in both Text Faithfulness and Visual Appeal. The reasons for this improvement can be summarized as follows: 1) Stronger foundation model: Layout-based models, such as LLM Blueprint, rely on Stable Diffusion 1.5 as the foundation model, which is less effective in text alignment and visual appeal compared to our ParaDiffusion. 2) Limitations of box-conditioned generation: Box-conditioned layouts are not always optimal for complex text alignment. As shown in Figure 12(c) and Figure 10, box-in-box scenarios frequently arise, making precise alignment challenging. Specifically, we summarize the following drawbacks of layout-based models: Overlap leads to misalignment in object count: When two identical objects overlap excessively, it often results in generating only one object. For example, two buildings may be represented as just one due to excessive overlap. Overlap causes object loss: When one object is inside another, it becomes difficult to generate the inner object. For instance, in the case of ’people in the boat,’ the person may fail to be generated. Incorrect relationships between objects: When two objects overlap, it can create issues in generating accurate object interactions. For example, ’a Samoyed lying on stones’ may be incorrectly generated as ’a Samoyed standing on stones. Therefore, current box-conditioned image generation models struggle to handle such overlapping relationships effectively. Additionally, as shown in Figure 10, layout-based models perform poorly in visual appeal. Many cases cannot be used, especially in LLM-grounded diffusion models, where a significant number of images are of poor quality.
4.2.5 Evaluation for The Detailed Concepts
In addition to exploring the overall assessment of Text Faithfulness, we also evaluated three more granular aspects: Objects, Attributes, and Locations, as shown in Figure 4. $50$ samples were randomly selected from the ParaPrompts 400 dataset, and we prompted GPT-4 to extract the Objects, Attributes, and Locations from each image description, as shown in Figure 13. This allowed us to obtain the number, names, attributes, and locations of objects in each prompt description. An evaluator was then tasked with providing a distribution score for each aspect, on a scale from $1$ to $5$, with $5$ being the highest. The results reveal that ParaDiffusion outperforms LLM Blueprint [8] in terms of alignment with Objects and Attributes, while it performs slightly worse in terms of Location alignment. The primary reason for this outcome is similar to previous analyses: layout-based image generation models, which use box-conditioned image generation (Paint by Example [48], BoxDiff [21]), achieve precise alignment for locations. However, these models also have significant limitations, such as issues with ”box-in-box” scenarios, which prevent them from effectively generating and fitting all objects and attributes. The corresponding object, attribute, and location annotations can be found at ParaPrompts-50-Concepts.
Additionally, to delve deeper into the comprehension and alignment of nuanced concepts of ParaDiffusion, along with the precision of object and attribute alignment, Figure 14 illustrates that our model adeptly tracks the semantic progression of the prompt. By sequentially incorporating pertinent object and attribute elements, ParaDiffusion demonstrates its ability to efficiently align with the intricacies of the textual instructions provided.
Table 10: Performance Comparison at Different Stages on ParaPrompts-400.
@@ 4.2.5 Evaluation for The Detailed Concepts
Pre-Train, Stage-1 - - - -
Stage-2 $vs.$ Stage-1 (w ParaImage-B) 82.1 17.9 76.9 32.1
Stage-3 $vs.$ Stage-2 (w ParaImage-S) 86.2 13.8 53.1 46.9
Table 11: Evaluation for Generated Caption. We used different model to generate captions for 500 images, scoring them from three perspectives on a scale of $1$ to $5$, with $5$ being the best.
Model Description Completeness Relevance Fantasy Ave. Score
CogVLM [44] 4.454 4.297 4.162 4.304
LLaVA 1.5 [26] 3.902 4.236 4.212 4.117
MiniGPT-4 [51] 3.744 3.523 3.233 3.500
@@ 4.3 Ablation Study
4.3.1 Effect of Language Model Adaptation.
@@
Table 6 presents the ablation study for LLM Adaptation. ‘Base’ denotes directly performing paragraph-image alignment learning (Stage-2) with ParaImage-Big without fine-tuning Llama V2 using LoRA. Compared to the base model, our LLM adaptation with LoRA demonstrates an improvement of nearly $5\%$ in human voting rates. Another insight reveals that during the process of increasing trainable parameters from $4.2$ million to $67.1$ million, the performance appears to gain consistently. Therefore, in our final configuration, we randomly select 16.7 million trainable parameters as the ultimate setting. Figure 12 (b) presents the visual comparisons for the performance from LLM adaptation. $r$ and $\alpha$ are two key hyperparameters, representing LoRA Rank and LoRA Alpha, respectively. Table 7 provides their detailed definitions and explanations. The performance comparison for different hyperparameter settings is presented in Table 6, where the optimal configuration is identified. The results indicate that when $r$=64, $\alpha$=32, the model achieves the best performance.
4.3.2 Effect of Caption Length.
Table 5 presents the ablation study of caption length for the proposed dataset, ParaImage. We control the same other variables, such as the size of training data and the images, and only change the corresponding caption length (words) through prompt CogVLM [44]. The specific prompt is as follows: ‘Provide a detailed description of the image using approximately $n$ words.’ It can be observed that as the length of the caption increases, the performance of the generated images in terms of text faithfulness also gradually increases. The data with average captions of around 150 words achieved the highest proportions of human voting, with $40.6\%$ and $31.2\%$ for ParaImage-Big and ParaImage-Small, respectively. It appears that an average of 150 words per caption still has not reached the peak performance. However, captioning capability of CogVLM is limited, unable to capture more granular descriptive information (exceeding average 150 words for the whole dataset).
4.3.3 Effect of Image Quality.
Table 8 presents the ablation for image quality of ParaImage. To ensure a fair comparison, for all datasets, we randomly sample $3$k data from the original dataset as the training set for training. It can be observed that the proposed ParaImage-Small exhibits the best performance, the models trained on this data achieving the highest human voting for visual appeal, at $54.7\%$. We believe this is reasonable, as the 3k images of ParaImage-Small were manually selected by annotators from a pool of 650k high-quality images of LAION-Aesthetics.
4.3.4 Effect of Data Size.
Table 9 presents the ablation study on the impact of data size. As the data size increases, performance of the model improves significantly across both automated metrics and human voting evaluations. Specifically, the model trained on 3.3 million ParaImage-Big data surpasses the one trained on 1.3 million, achieving a 17.5% improvement in human voting. Similarly, in the automated metric FID-30K, the score improves notably from 13.45 to 9.64.
4.3.5 Performance Comparison for Different Stages.
@@
Table 10 presents performance comparison at different stages. ParaImage-Big brings significant improvements, with approximately a $50\%$ increase in human voting rates for both visual appeal and text faithfulness simultaneously. The smaller-scale, and high-quality set of 3k images in ParaImage-Small further enhances aesthetic aspects, achieving a remarkable around $70\%$ increase in human preference rates compared to that of Stage 2. Figure 12 (a) presents a visual comparison, vividly illustrating the gains in visual appeal. In this case, the prompt is ‘An exquisite sculpture in the center of a square on a rainy day.’
4.3.6 Quality Evaluation for Generative Captions of ParaImage-Big
As demonstrated in Section 3.2.1, we utilized a substantial amount of synthetic caption data. However, the quality of synthetic captions still leaves room for improvement and does not match the precision of manually annotated captions (ParaImage-Small). Here, we assessed the quality of synthetic captions. As indicated in Table 11, we primarily selected three mainstream vision-language models (CogVLM, LLaVA, MiniGPT-4) to generate corresponding captions, which were then manually evaluated for quality. We mainly evaluated the attributes of the captions (Description Completeness, Relevance, Fantasy) from three aspects. In this evaluation, three assessors were invited. For each generated caption, they were scored on a scale from 1 to 5, and we took the average of their scores as the final score for that caption. Among these three models, CogVLM performed the best, achieving a score of $4.304$.
4.3.7 Potential Biases in Human Evaluation
Human evaluation is crucial for assessing generative models, but it is important to acknowledge several potential biases that may influence the results:
  • Evaluator Bias. Different evaluators may have varying personal preferences, cultural backgrounds, or levels of expertise, which could affect their ratings. To reduce this bias, one effective approach is to ensure diversity in our evaluator pool and provide clear, standardized rating criteria.
  • Rule Bias. Evaluators may interpret the rating scale differently, leading to inconsistencies in scores. To address this, one effective approach is to provide detailed guidelines and examples for each score level, along with a calibration process to align evaluators’ understanding of the scale. For example, in Sections 7.5 and 3.2.2, we implemented constraints on caption labeling and image selection to enhance objectivity and reduce individual subjectivity.
  • Task Familiarity Bias: Familiarity with the task can affect how evaluators assess the generated images. Experts might focus on technical details, while novices may prioritize overall aesthetics.
While human evaluations are essential, they are subject to various biases. Therefore, in conducting our user study, we carefully considered these biases and established comprehensive evaluation rules to minimize subjective influences. It is also important to expand the evaluator pool to ensure diversity across different backgrounds, aiming to make the evaluation results more objective and fair.
5 Research Data Policy and Data Availability Statements
  • The datasets used during and/or analysed during the current study are available in the related public dataset repositories, according to Sec. 3.2. Or more detailed information can be found at our repository https://github.com/weijiawu/ParaDiffusion.
  • For the datasets introduced in this paper (i.e. ParaImage), we have open-sourced the training and testing sets, as well as the corresponding methods for data collection, screening, annotation, and evaluation. Please refer to the respective dataset GitHub link for specific details.
@@ 7.1 More Visualizations
Figure 18 and Figure 15 provides more visualizations of ParaDiffusion for human-centric and scenery-centric domains, respectively. The visualization reveals that our ParaDiffusion can generate intricate and realistic composite images of individuals as well as picturesque scenes. Moreover, it is noteworthy that images generated through paragraph-image generation exhibit a compelling narrative quality, enriched with profound semantics. The generated images consistently feature intricate object details and demonstrate effective multi-object control.
@@ 7.3 Experiments on 1,600 prompts of PartiPrompts
We also provides the related experiment results on
PartiPrompts-1600, as shown in Figure 17. It can be observed that our model also achieved outstanding performance in Text Faithfulness, with a $27.3\%$ human voting rate, significantly outperforming previous models such as SD XL and DeepFloyd-IF. Additionally, our model demonstrated a competitive advantage in visual appeal, surpassing SD XL and DeepFloyd-IF, and approaching the performance of contemporaneous work PIXART-$\alpha$. A notable observation in ParaPrompts is the high proportion of ’Tie’ votes in human evaluations, especially for Text Faithfulness, with a voting rate of up to $34\%$. This is attributed to the presence of numerous simple or abstract prompts in ParaPrompts-1600, making it challenging to provide precise voting results, such as for prompts like ‘happiness’ or ‘emotion.’